OpenStudy (anonymous):

Okay so I'm completely lost.... Does anybody understand operations of on complex numbers? Basically problems like (0-2i)-(8-6i)??? Thanks!

OpenStudy (campbell_st):

well its just like algebra... its about collecting like terms. so remove the backets and you ge 0 - 2i - 8 + 6i collect the like terms ( 0 - 8) + (-2i + 6i) hopefully this allows you to find the answer.

OpenStudy (anonymous):

Wouldn't removing the brackets put it to 0-2i-8-6i???

OpenStudy (campbell_st):

well you need to distribute -1 so in the 2nd set of brackets its -1(8 - 6i) is -1 x 8 = -8 -1 x -6i = 6i hope that helps

OpenStudy (yttrium):

@xXxBambyGirlxXx ( 0 - 2i) - (8 - 6i ) is like ( 0 - 2i) -1 (8 - 6i ) --> so there you will need to substitute -1 then it will become (0 - 2i) + (-8+6i), right? then removing the parentheses will give you 0-2i-8+6i which is equal to 4i-8 or is also similar to 4 (i - 2)

OpenStudy (anonymous):

I'm lost.... /:

OpenStudy (campbell_st):

ok... so look at the 2nd set of brackets is \[-1 \times (8 - 6i)\] can you distribute the -1...?

OpenStudy (anonymous):

By making -8+6i?

OpenStudy (campbell_st):

thats correct.... and the 1st set of brackets is \[1 \times (0 - 2i)\] can you distribute that..?

OpenStudy (anonymous):

It would just be 0-2i correct?

OpenStudy (anonymous):

JUST TREAT I AS ANY VARIABLE AND JUST OPEN THE BRACKETS AND SOLVE JUST REMEMBER THAT REAL AND IMAGINARY PARTS DO NOT ADD OR SUBTRACT WITH EACH OTHER

OpenStudy (anonymous):

ANSWER WOULD BE -8+4i

OpenStudy (campbell_st):

thats correct... so after distributing it looks like 0 - 2i - 8 + 6i now collect the like terms...

OpenStudy (anonymous):

-8+4i!(:

OpenStudy (campbell_st):

thats it.... just like those basic algebra classes where you collected like terms... well done

OpenStudy (anonymous):

Could you walk me through the next one?..

OpenStudy (campbell_st):

ok... if you post it in the normal way I'll look at it...

OpenStudy (anonymous):

Multiply.. (6+4i)(5-9i)

OpenStudy (campbell_st):

ok... so I'll do this the long way by splitting the 1st brackets... \[6(5 - 9i) + 4i(5 - 9i)\] can you distribute them...?

OpenStudy (anonymous):

Okay okay so... 30-54i+20i-36i??

OpenStudy (campbell_st):

thats great... but the last term is 36i^2... does that make sense...?

OpenStudy (campbell_st):

so the problem is now 30 - 54i + 20i - 36i^2

OpenStudy (anonymous):

Ohhh yeah I thought so!

OpenStudy (campbell_st):

ok... cool... now do you know about i^2 = -1...?

OpenStudy (anonymous):

Okay so combine like terms... No so it would basically be 30-54i+20i-36??

OpenStudy (campbell_st):

and substituting... before simplifying you get \[30 - 54i + 20i - 36 \times (-1) = 30 -54i + 20i + 36\] does that make sense...?

OpenStudy (campbell_st):

now you can combine the like terms...

OpenStudy (anonymous):

Okay okay! so 66-34i??

OpenStudy (campbell_st):

thats it... well done

OpenStudy (anonymous):

So would (5-3i)(2-8i) come out to be: -14-46i?

OpenStudy (campbell_st):

yes... that looks great... well done took me a minute to do it in my head

OpenStudy (anonymous):

Haha now addition... Help me one last time?

OpenStudy (campbell_st):

ok... as long as it's quick... post the question and you're solution

OpenStudy (anonymous):

(4-2i)+(12+7i) I'm not sure where to start on this one /:

OpenStudy (anonymous):

-8+9i?

OpenStudy (campbell_st):

ok... so with both sets of brackets its \[1 \times (4 - 2i) + 1 \times (12 + 7i) = 4 - 2i + 12 + 7i\] collect the like terms...

OpenStudy (campbell_st):

so you get 4 + 12 +7i - 2i = can you finish..?

OpenStudy (anonymous):

16+5i?

OpenStudy (campbell_st):

thats it.... well done